# Variational Learning Finds Flatter Solutions at the Edge of Stability

### v2

@@ 3 Stability Threshold for A Simple Case of Variational Learning
Figure 2: (a) Theoretical stability threshold $2/\rho\cdot\text{VF}(z)$ as a function of $z=N_{s}/\sigma^{2}$. (b) Empirical descent probability $\mathbb{P}(\ell(m_{t+1})-\ell(m_{t})<0)$ over the grid $(1/\sigma^{2},\lambda)$ on the loss $f(m)=(\lambda/2)\cdot m^{2}$, shown as a grayscale heatmap; white indicates high descent probability of loss decrease. The curve $\lambda=2/\rho\cdot\text{VF}(z)$ denotes the stability threshold. (c) Stability boundary over $(N_{s},\sigma^{2})$, with instability observed at low $N_{s}$ and high $\sigma^{2}$.
So far, we show that the stability threshold for Variational GD is smaller than that of GD, but how can we control it? We now show some visualization which show the explicit behavior of the threshold as a function of $N_{s}$ and $\sigma_{i}^{2}$. We consider a 1D quadratic loss $\ell(m)=(\lambda/2)\cdot m^{2}$, for which GD with step-size $\rho$ will decrease the loss in one step when $\lambda<2/\rho$. To validate Theorem 5, we run Variational GD for several combination of losses with different $\lambda$ and posteriors with different variance $\sigma^{2}$. We fixed the step-size $\rho$ and number of samples $N_{s}$. For each combination $(1/\sigma^{2},\lambda)$, we compute the empirical descent probability $\mathbb{P}(\ell(m^{\epsilon}_{t+1})-\ell(m_{t})<0)$ by running one iteration of the update across 10 realizations of $\epsilon$ and recording the fraction of runs where the loss decreased.
The frequency is visualized as a heatmap in Figure 2(b), where white indicates high descent probability and black indicates low. The red curve denotes the stability threshold from Theorem 5 and we observe that the transition from low to high descent probability aligns with the stability threshold. This experiment empirically confirms Theorem 5 and Lemma 3, showing that when the curvature $\lambda$ lies below the stability threshold (solid red curve), the loss decreases with high probability. Additional stability curves for varying posterior sample sizes $N_{s}$ are provided in Appendix C.
The stability threshold can be controlled by varying both the variance $\sigma^{2}$ and the number of samples $N_{s}$. In Figure 2(c), we show the linear relationship between the variance $\sigma^{2}$ and samples $N_{s}$ that controls the stability threshold. Values of $(N_{s},\sigma^{2})$ lying above the dashed red line leads to unstable dynamics in a quadratic. In the next section, we analyze the role of this instability mechanism on a quadratic loss in regulating the loss Hessian in deep neural networks.
@@ 3.1 Nature of Stability in GD and VGD
A fixed point $\boldsymbol{\theta}^{*}$ of an iterative algorithm is said to be asymptotically stable if it is Lyapunov stable and there also exists a $\delta>0$ such that if the algorithm starts within this $\delta$-neighborhood of the fixed point, the iterates will converge to the fixed point. Formally, if $\|\boldsymbol{\theta}_{0}-\boldsymbol{\theta}^{*}\|<\delta$, then:
@@
GD on a quadratic is asymptotically stable because the condition learning-rate $\rho<2/\lambda_{1}$ ensures that the iteration matrix $(\mathbf{I}-\rho\mathbf{Q})$ has all eigenvalues less than $1$, guaranteeing convergence of the iterates to its fixed point. For VGD, the iterates are Stochastically Stable which is defined as follows
@@
Figure 3: Smaller sharpness corresponds to higher test accuracy for various network architectures trained on the CIFAR-10 dataset. Subfigures show the results for (a) ViT, (b) MLP, and (c) ResNet-20. Full Batch GD with Variational GD consistently achieves lower sharpness and better test accuracy.
@@ 4 Experiments on Deep Neural Networks
Similarly to the GD case, we expect the stability threshold for the quadratic to serve as an EoS limit. That is, we expect that, when using variational learning for deep neural networks, the sharpness should hover around the stability limit. By controlling the covariance shape and number of Monte Carlo sample we can stere optimization into regions where sharpness is much lower than GD. The following hypothesis formally states this intuition which we verify through extensive experiments.
@@
For a twice-differentiable, non-convex loss $\ell(\mathbf{m})$ optimized using the update rule (5) with Gaussian noise $\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ with variance $\sigma^{2}$, the top Hessian eigenvalue $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ hovers around the stability threshold
@@
Figure 4: Normalized Sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}/(2/\rho)$ hovers about the Variational Factor in MLP.
@@
Perturbation effect of Gaussian variance: In Figure 5, we present experiments on weight-perturbed Gradient Descent (GD) for a classification task using a multi-layer perceptron (MLP), where we plot the sharpness of the final iterate as a function of the variance $\sigma^{2}$ of an isotropic Gaussian. For three different learning rates, $\rho=0.05$, $0.1$, and $0.2$, we run variational GD with different $\sigma^{2}$. Across all learning rates $\rho$ and numbers of posterior samples $N_{s}$, we observe that larger variance consistently leads to lower sharpness. This exactly aligns with Hypothesis 1, larger perturbations reduce the stability threshold, promoting escape from sharper regions of the loss landscape.
@@
(a) $402/\rho=$ 40
(b) $202/\rho=$ 20
(c) $102/\rho=$ 10
Figure 5: Sharpness of the final iterate for a MLP network trained on CIFAR-10 for variational GD achieved across various learning-rate $\rho$, noise covariance $\boldsymbol{\Sigma}$ and posterior samples $N_{s}$. Higher variance and smaller samples lead to lower sharpness. $X$ denotes divergence.
Comparison with regularization effect of SGD: Our work differs fundamentally from studies on the regularization effects of mini-batch noise in SGD, such as Wu et al. (2022); Zhu et al. (2019); Wu et al. (2018); Ibayashi and Imaizumi (2023); Mulayoff and Michaeli (2024), which analyze gradient noise and its role in promoting flatter solutions. In contrast, variational GD introduces noise directly in the weights, inducing a structured and anisotropic form of gradient noise shaped by the curvature of the loss landscape, something not captured by existing SGD analyses. While most SGD-based studies assume Gaussian gradient noise, despite empirical evidence of heavy-tailed behavior (Gurbuzbalaban et al., 2021; Nguyen et al., 2019), we show that Gaussian perturbations in weights lead to Gaussian-distributed gradients, that is, $\hat{\mathbf{g}}\sim\mathcal{N}(\nabla\ell(\mathbf{m}_{t}),\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q})$, where the covariance is amplified along sharper directions. This formulation requires no assumptions on the shape of gradient noise (unlike, e.g., Lee and Jang (2023)). To complement our work, we further perform an empirical study using weight perturbations drawn from a heavy-tailed distribution in deep neural networks.
@@ 4.1 Edge of Stability under Heavy-Tailed Noise
In weight-perturbed gradient descent (GD), the stability threshold depends not only on the perturbation covariance and sample size, but also on the posterior’s tail behavior. We empirically investigate this by drawing weight perturbations from a Student-t distribution with degrees of freedom $\alpha$, where smaller $\alpha$ induces heavier tails and larger deviations in the perturbed gradients. The distribution becomes heavier-tailed as $\alpha$ decreases, with the Gaussian recovered as $\alpha\to\infty$. In Figure 6, we train an MLP on CIFAR-10 and observe that heavier-tailed perturbations yield lower sharpness and better test accuracy. Similar trends are confirmed for ResNet-20 and ViT in the Appendix F. While a theoretical analysis under heavy-tailed noise is challenging, our results highlight that the posterior shape plays a critical role in generalization.
Figure 6: Sharpness dynamics with noise injection from heavy-tailed student-t posterior distribution parameterized by $\alpha$. Perturbation sampled from a heavier tail (smaller $\alpha$) leads to smaller sharpness.
4.2 Adaptive Edge of Stability in Variational Online Newton
To solve the variational objective (1), one option is to use natural-gradient descent over the parameters of the Gaussian posterior $q_{t}=\mathcal{N}(\mathbf{m}_{t},\mathbf{P}_{t}^{-1})$. This leads to the Variational Online Newton (VON) update rule (Khan and Rue, 2023, Eq. 12),
@@ 4.2 Adaptive Edge of Stability in Variational Online Newton
The posterior covariance $\mathbf{P}_{t}^{-1}$ is learned online using second-order curvature estimate. The variational GD in Eq. (5) corresponds to the special case where $\mathbf{P}_{t}$ is fixed across iterations. Adaptive optimizers such as Adam, RMSProp, and Adadelta can be seen as special cases of VON, as shown in (Khan and Rue, 2023, Section 4.2). Adaptive gradient methods differ from standard gradient descent as the former can adapt their preconditioner $\mathbf{P}_{t}$ and move into high curvature regions. For example in Adam, the preconditioner $\mathbf{P}_{t}$ is updated in an exponential moving average (EMA) fashion with parameter $\beta_{2}$:
Here, all operations such as squaring or division of vectors are performed elementwise. Recent work by Cohen et al. (2022) studies the dynamics of adaptive gradient methods, introducing the concept of Adaptive Edge of Stability (AEoS). They show that instead of sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$, a modified quantity, termed preconditioned sharpness $\|\mathbf{P}^{-1}_{t}\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ hovers about $2/\rho$.
Figure 7: Pre-conditioned sharpness of Adam and IVON with various temperatures ($\tau$). Smaller temperatures which shrinks the posterior reaches larger preconditioned sharpness.
Here, we study the AEoS for a recent large-scale implementation of VON by Shen et al. (2024) called IVON (Improved Variational Online Newton). There, the update for the preconditioner is approximated using Stein’s identity to estimate the Hessian from $N_{s}$ posterior samples:
@@ 4.3 Effect of Batch Size
Figure 8: Minimizing Variational Objective depends on the choice of hyperparameter and its implicit regularization effect.
@@ Appendix A Related Works
Given a prior and a likelihood, variational inference approximates the posterior by optimizing the evidence lower bound within a family of posterior distribution. The ELBO is optimized by natural gradient descent Amari [1998], in the hope that steepest descent induced by the KL divergence is a better metric to compare probability distributions, this gives to rise of a class of algorithms called Natural Gradient Variational Inference (NGVI). Several works have attempted to apply NGVI in optimizing deep networks, albeit with higher costs or tricky implementations (such as) Khan et al. [2018], Osawa et al. [2019], Lin et al. [2020]. However, recently Shen et al. [2024] provided an improved version of variational Online Newton that largely scales and obtains state of the art accuracy and uncertainty at identical cost as Adam. A step of VL includes weight perturbation based on noise injection from the posterior distribution which has similarities to noise injection.
@@
Appendix B Proof of Theorem-1
@@
1 $1~{}-\;2\,\exp\!\Bigl{(}-\,d_{1}\,\min\Bigl{\{}\tfrac{\delta^{2}\,N_{s}^{2}}{d_{2}},\,\tfrac{\delta\,N_{s}}{d_{2}}\Bigr{\}}\Bigr{)},$
we have $\Delta\ell<0$. Thus, if the expected descent step is at most $-\delta$, then $\Delta\ell$ is negative with high probability. ∎
@@ Appendix C Role of posterior samples on a quadratic
In Figure 12, we visualize the iterate dynamics on a quadratic loss under varying posterior variances and sample sizes. As the plots show, larger variance and smaller $N_{s}$ lead to increased instability in the iterates. This instability mechanism can facilitate generalization in deep neural networks by enabling escape from sharper minima.
@@ Appendix D Loss Smoothing Beyond Local Quadratic Approximation
For example, let’s take the example of a quartic function $\ell(\theta)=(\theta^{2}-1)^{2}$ where minima is at $\theta^{*}=\pm 1$ and ′′$\ell^{{}^{\prime\prime}}(\theta^{*})=12{\theta^{*}}^{2}-4=8$. Now, the smoothed loss $\ell_{conv}(m)=\mathbb{E}_{q}\ell(\theta)=\mathbb{E}_{q}(\theta^{2}-1)^{2}=m^{4}+m^{2}(6\sigma^{2}-2)+(3\sigma^{4}-2\sigma^{2}+1)$. The new loss has a global minima at $m^{*}=\pm\sqrt{1-3\sigma^{2}}$ (only if $\sigma<\sqrt{\frac{1}{3}}$). Comparing the curvature of both the losses at the global minima, we have:
′′$\ell^{{}^{\prime\prime}}(\theta)=12{\theta}^{2}-4=8$
@@
[…]
@@
Note that $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is sub-Gaussian, and $\epsilon_{i}^{2}$ is sub-exponential. Thus $Y_{i}=6\,\epsilon_{i}^{2}+4\,\epsilon_{i}$ is a linear combination of a sub-exponential and a sub-Gaussian variable, which remains sub-exponential. Shifting by the constant  $-\,6\,\sigma^{2}$ does not affect sub-exponential parameters. Therefore, each $Y_{i}$ is sub-exponential.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$N_{s}$ $\sigma^{2}=0.1$ $\sigma^{2}=0.5$ $\sigma^{2}=1.0$ $\sigma^{2}=2.0$
1000 Refer to caption Refer to caption Refer to caption Refer to caption
50 Refer to caption Refer to caption Refer to caption Refer to caption
10 Refer to caption Refer to caption Refer to caption Refer to caption
5 Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Instability of variational GD on a quadratic with varying posterior samples $N_{s}$ and varying posterior variance $\sigma^{2}$.
@@ 5 Refer to caption Refer to caption Refer to caption Refer to caption
[…]

### v3

@@ 3 Stability Threshold for a Simple Case of Variational Learning
Figure 3: (a) Solid black curve shows the theoretical stability-threshold of VGD as a function of $N_{s}/\sigma^{2}$. The curve is clearly lower than the stability threshold of GD, shown with the horizontal dashed, gray line. (b) Empirical verification on a scalar quadratic problem with curvature $\lambda$ where we plot the empirically computed probability of descent for VGD runs with different values of $\sigma^{2}$. We show a heatmap for $(\lambda,1/\sigma^{2})$ values where lighter colors indicate higher probability of descent. We overlay the heatmap with the stability thresholds of VGD (red solid curve), clearly showing that theoretical limit shown in (6) matches the empirical probability. (c) The figure further includes $N_{s}$ and marks the region where a pair $(N_{s},\sigma^{2})$ will either lead to descent or not (marked with ‘stable’ and ‘unstable’ respectively).
So far, we show that the stability threshold for Variational GD is smaller than that of GD, but can we also ensure it to be smaller in practice? The answer is yes, and it can be done by controlling $\sigma^{2}$ and $N_{s}$. We will now demonstrate this on a 1D quadratic loss $\ell(m)=\mbox{$\frac{1}{2}$}\lambda m^{2}$. For such cases, GD with step-size $\rho$ descends whenever the curvature $\lambda$ is bounded by $2/\rho$, but we can show that VGD descends only for lower curvature values. To show this, we run VGD for many $(\lambda,\sigma)$ pairs. For each run, we use 10 random realizations of $\epsilon$, then record how often the loss decreases, and finally compute the approximate descent probability $\mathbb{P}\big(\ell(m_{t+1}^{\epsilon})<\ell(m_{t})\big)$.
Figure 3(b) shows this probability as a heatmap where lighter color indicate a higher probability of descent; a white pixel indicates a probability of 1 and a black one indicates a probability of 0. We overlay this with the solid red curve showing the theoretical stability-threshold of Variational GD as dictated by (6), that is, $\lambda=(2/\rho)\,\mathrm{VF}(N_{s}/\sigma^{2})$. The theoretical curve closely matches the transition boundary where probabilities transition from 1 to 0. The figure clearly shows that by increasing $\sigma^{2}$ we can reduce the stability threshold in practice too. Figure 3(c) further illustrates the effect of changing $N_{s}$, showing the regions where a $(N_{s},\sigma^{2})$ pair lead to descent (marked as ‘stable’) or otherwise (marked with ‘unstable’). As expected, the relationship is linear and the same effect is obtained by either increasing $N_{s}$ or decreasing $\sigma^{2}$.
@@ 3.1 Nature of Stability in GD and VGD
Let $\boldsymbol{\theta}^{*}$ be the minimum. An iterate $\boldsymbol{\theta}_{k}$ is asymptotic stable if it is Lyapunov stable and there also exists a $\delta>0$ such that if the algorithm starts within a $\delta$-neighborhood of the fixed point, the iterate will converge to the fixed point. Formally, if $\|\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}\|<\delta$ for a finite $k$, then:
@@
GD on a quadratic is asymptotically stable because the condition learning-rate $\rho<2/\lambda_{1}$ ensures that the iteration matrix $(\mathbf{I}-\rho\mathbf{Q})$ has all eigenvalues less than $1$, guaranteeing convergence of the iterates to its fixed point. Analogously VGD, the iterates are Stochastically Stable which is defined as follows
@@
3.2 Comparison with Regularization Effect of SGD
Our work differs fundamentally from studies on the regularization effects of mini-batch noise in SGD, such as Wu et al. (2022); Zhu et al. (2019); Wu et al. (2018); Ibayashi and Imaizumi (2023); Mulayoff and Michaeli (2024), which analyze gradient noise and its role in promoting flatter solutions. In contrast, VGD introduces noise directly in the weights, inducing a structured and anisotropic form of gradient noise shaped by the curvature of the loss landscape, something not captured by existing SGD analyses. While most SGD-based studies assume Gaussian gradient noise, despite empirical evidence of heavy-tailed behavior (Gurbuzbalaban et al., 2021; Nguyen et al., 2019), we show that for the quadratic loss, Gaussian perturbations in weights lead to Gaussian-distributed gradients, that is, $\hat{\mathbf{g}}\sim\mathcal{N}(\nabla\ell(\mathbf{m}_{t}),\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q})$, where the covariance is amplified along sharper directions. This formulation requires no assumptions on the shape of gradient noise (unlike, e.g., Lee and Jang (2023)). To complement our work, we further perform an empirical study using weight perturbations drawn from a heavy-tailed distribution in deep neural networks.
@@ 3.2 Comparison with Regularization Effect of SGD
Figure 4: Smaller sharpness corresponds to higher test accuracy for network architectures trained on CIFAR-10. Panels show (a) ViT, (b) MLP, and (c) ResNet-20. Full batch GD with Variational GD achieves lower sharpness and better test accuracy.
@@ 4 Experiments on Deep Neural Networks
Similarly to the GD case, we expect the stability threshold for the quadratic to serve as an EoS limit. That is, we expect that, when using variational learning for deep neural networks, the sharpness should hover around the stability limit. By controlling the covariance shape and number of Monte Carlo samples we can stere optimization into regions where sharpness is much lower than GD. The following hypothesis formally states this intuition which we verify through extensive experiments.
@@
For a twice-differentiable loss $\ell(\mathbf{m})$ optimized using the update rule (5), the top Hessian eigenvalue $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ hovers around the stability threshold,
@@
Figure 5: Normalized Sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}/(2/\rho)$ hovers around the Variational Factor in MLP.
@@
Perturbation effect of Gaussian variance: In Figure 6, we present experiments on VGD for a classification task using a multi-layer perceptron (MLP), where we plot the sharpness of the final iterate as a function of the variance $\sigma^{2}$ of an isotropic Gaussian. For three different learning rates, $\rho=0.05$, $0.1$, and $0.2$, we run VGD with different $\sigma^{2}$. Across all learning rates $\rho$ and numbers of posterior samples $N_{s}$, we observe that larger variance consistently leads to lower sharpness. This exactly aligns with Hypothesis 1, larger perturbations reduce the stability threshold, promoting escape from sharper regions of the loss landscape.
@@
(a) $2/\rho=40$
(b) $2/\rho=20$
(c) $2/\rho=10$
Figure 6: Sharpness of the final iterate for an MLP trained on CIFAR-10 for variational GD across learning rate $\rho$, noise covariance $\boldsymbol{\Sigma}$, and posterior samples $N_{s}$. Higher variance and smaller samples lead to lower sharpness. For panel (c), training did not converge for variance values 2 and 5, therefore not shown in the plot.
@@ 4.1 Edge of Stability under Heavy-Tailed Noise
In variational gradient descent (VGD), the stability threshold depends not only on the perturbation covariance and sample size, but also on the posterior’s tail behavior. We empirically investigate this by drawing weight perturbations from a Student-t distribution with degrees of freedom $\alpha$, where smaller $\alpha$ induces heavier tails and larger deviations in the perturbed gradients. The distribution becomes heavier-tailed as $\alpha$ decreases, with the Gaussian recovered as $\alpha\to\infty$. In Figure 7, we train an MLP on CIFAR-10 and observe that heavier-tailed perturbations yield lower sharpness and better test accuracy. Similar trends are confirmed for ResNet-20 and ViT in Appendix F. While a theoretical analysis under heavy-tailed noise is challenging, our results highlight that the posterior shape plays a critical role in generalization.
Figure 7: Sharpness dynamics with noise injection from a heavy tailed Student t posterior parameterized by $\alpha$. Perturbations from heavier tails (smaller $\alpha$) lead to smaller sharpness.
4.2 Adaptive Edge of Stability and Variational Online Newton Methods
Recent work by Cohen et al. (2022) studies the dynamics of adaptive gradient methods, introducing the concept of Adaptive Edge of Stability (AEoS). In this section, we will show that a similar phenomenon happens in natural-gradient variational learning.
Cohen et al. (2022) show that for adaptive gradient methods instead of sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$, a modified quantity, termed preconditioned sharpness $\|\text{diag}(\mathbf{p}_{t})^{-1}\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ hovers around $2/\rho$, where $\mathbf{p}_{t}$ is a preconditoner. Adaptive gradient methods differ from standard gradient descent as the former can adapt their preconditioner $\mathbf{P}_{t}$ and move into high curvature regions. For example in Adam, a preconditioner $\mathbf{p}_{t}$ is updated in an exponential moving average (EMA) fashion with parameter $\beta_{2}$:
Here, all operations such as squaring or division of vectors are performed elementwise.
Natural-gradient VL methods which learn a complete Gaussian posterior $q_{t}=\mathcal{N}(\mathbf{m}_{t},\mathbf{P}_{t}^{-1})$ take a similar form to the above adaptive gradient methods. An instance of this is the Variational Online Newton (VON) update rule (Khan and Rue, 2023, Eq. 12),
@@ 4.2 Adaptive Edge of Stability and Variational Online Newton Methods
Here, the posterior covariance $\mathbf{P}_{t}^{-1}$ is learned using the loss Hessians. The variational GD in Eq. (5) corresponds to the special case where $\mathbf{P}_{t}$ is fixed across iterations. Adaptive optimizers such as Adam, RMSProp, and Adadelta can be seen as special cases of VON, as shown in (Khan and Rue, 2023, Section 4.2).
Figure 8: Preconditioned sharpness for Adam and IVON across temperatures $\tau$. Smaller $\tau$ shrinks the posterior and yields larger preconditioned sharpness.
Here, we study the AEoS for a recent large-scale implementation of Equation˜11 by Shen et al. (2024) called IVON (Improved Variational Online Newton). There, the update for the preconditioner is approximated using Stein’s identity to estimate the Hessian from $N_{s}$ samples:
@@ 4.3 Effect of Batch Size
Figure 9: Minimizing the variational objective depends on the choice of hyperparameter and its implicit regularization effect, visible in both the objective value and sharpness reduction.
@@ Acknowledgements
  • Related Works ........................................................................................................................................................................A
  • Proof of Theorem 1 ........................................................................................................................................................................B
  • Role of posterior samples on a quadratic ........................................................................................................................................................................C
  • Loss Smoothing Beyond Local Quadratic Approximation ........................................................................................................................................................................D
  • Discussion on stability and descent for Variational GD ........................................................................................................................................................................E
  • Additional experiments across diverse datasets ........................................................................................................................................................................F
@@ Appendix A Related Works
Given a prior and a likelihood, variational inference approximates the posterior by optimizing the evidence lower bound within a family of posterior distribution. The ELBO is optimized by natural gradient descent [Amari, 1998], in the hope that steepest descent induced by the KL divergence is a better metric to compare probability distributions, this gives to rise of a class of algorithms called Natural Gradient Variational Learning. Several works have attempted to apply this to deep learning [Khan et al., 2018, Osawa et al., 2019, Lin et al., 2020]. Recently, Shen et al. [2024] provide an improved version of varlational Online Newton that largely scales and obtains state of the art accuracy and uncertainty at identical cost as Adam. A step of VL includes weight perturbation based on noise injection from the posterior distribution which has similarities to noise injection.
@@
Appendix B Proof of Theorem 1
@@
For simplicity, we assumed to $\mathbf{Q}$ to be full rank. If $\mathbf{Q}$ has a non-trivial null-space that is $\lambda_{i}=0$ for some $d>i\geq k$, then we have $\mathbf{v}_{i}^{\top}\mathbf{g}=\mathbf{v}_{i}^{\top}\mathbf{Q}\mathbf{m}_{t}=\lambda_{i}\mathbf{v}_{i}^{\top}\mathbf{m}_{t}=0$. This means there is no component of the iterate in the null-space direction of $\mathbf{Q}$ and does not contribute to the descent objective.
@@
$1~-\;2\,\exp\!\Bigl(-\,d_{1}\,\min\Bigl\{\tfrac{\delta^{2}\,N_{s}^{2}}{d_{2}},\,\tfrac{\delta\,N_{s}}{d_{2}}\Bigr\}\Bigr),$
we have $\Delta\ell<0$. Thus, if the expected descent step is at most $\delta$, then $\Delta\ell$ is negative with high probability. ∎
@@ Appendix D Loss Smoothing Beyond Local Quadratic Approximation
For example, let’s take the example of a quartic function $\ell(\theta)=(\theta^{2}-1)^{2}$ where minima is at $\theta^{*}=\pm 1$ and $\ell^{{}^{\prime\prime}}(\theta^{*})=12{\theta^{*}}^{2}-4=8$. Now, the smoothed loss $\ell_{conv}(m)=\mathbb{E}_{q}\ell(\theta)=\mathbb{E}_{q}(\theta^{2}-1)^{2}=m^{4}+m^{2}(6\sigma^{2}-2)+(3\sigma^{4}-2\sigma^{2}+1)$. The new loss has a global minima at $m^{*}=\pm\sqrt{1-3\sigma^{2}}$ (only if $\sigma<\sqrt{\frac{1}{3}}$). Comparing the curvature of both the losses at the global minima, we have:
$\ell^{{}^{\prime\prime}}(\theta)=12{\theta}^{2}-4=8$
@@
$\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\;=\;\frac{2}{N_{s}}\sum_{i=1}^{N_{s}}\Bigl[\,(6\,\epsilon_{i}^{2}+4\,\epsilon_{i})\;-\;6\,\sigma^{2}\Bigr].$
@@
Note that $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is sub-Gaussian, and $\epsilon_{i}^{2}$ is sub-exponential. Thus $Y_{i}=6\,\epsilon_{i}^{2}+4\,\epsilon_{i}$ is a linear combination of a sub-exponential and a sub-Gaussian variable, which remains sub-exponential. Shifting by the constant  $6\,\sigma^{2}$ does not affect sub-exponential parameters. Therefore, each $Y_{i}$ is sub-exponential.
@@
$\Pr\!\Bigl(\,\Bigl|\tfrac{1}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}\Bigr|\,\geq\,t\Bigr)\;\;\leq\;\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{t^{2}}{v},\,\tfrac{t}{b}\Bigr\}\Bigr).$
@@
$\Bigl|\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\Bigr|\;=\;\frac{2}{N_{s}}\,\Bigl|\sum_{i=1}^{N_{s}}Y_{i}\Bigr|\;=\;2\,\Bigl|\tfrac{1}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}\Bigr|.$
@@
$\Pr\!\Bigl(\,\bigl|\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\bigr|\;\geq\;\delta\Bigr)\;=\;\Pr\!\Bigl(\,\Bigl|\tfrac{1}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}\Bigr|\;\geq\;\tfrac{\delta}{2}\Bigr)\;\;\leq\;\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{(\delta/2)^{2}}{v},\,\tfrac{\delta/2}{b}\Bigr\}\Bigr).$
@@
$\Pr\!\Bigl(\,\bigl|\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\bigr|\;\geq\;\delta\Bigr)\;\leq\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{\delta^{2}}{4c_{1}\sigma^{4}},\;\tfrac{\delta}{2c_{2}\sigma^{2}}\Bigr\}\Bigr).$
@@
$\Pr\!\Bigl(\,\bigl|f_{\mathrm{avg}}^{\prime\prime}(x)\;-\;f_{\mathrm{conv}}^{\prime\prime}(x)\bigr|\,\geq\,\delta\Bigr)\;\leq\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{\delta^{2}}{4\,c_{1}\,\sigma^{4}},\;\tfrac{\delta}{2\,c_{2}\,\sigma^{2}}\Bigr\}\Bigr).$